OpenStudy (anonymous):

f(x)=(e)^(-4x^(2)) need f of 1 and f of 2.

OpenStudy (anonymous):

just plgu in 1 for x and 2 for x

OpenStudy (anonymous):

f(x)=(e)^(-4x^(2)) f(1)=(e)^(-4(1)^(2)) f(2)=(e)^(-4(2)^(2))

OpenStudy (anonymous):

i meant f'(x) and f"(x)

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

lol f(x) means f of x lol umm use the produce rule again...we already showed you in a different problem

OpenStudy (anonymous):

i know. i keep getting the wrong answers though. i suck at this.

OpenStudy (anonymous):

actually you don't use the product rule because its e^

OpenStudy (anonymous):

and how would i use the product rule theres only an e raised to something

OpenStudy (anonymous):

you know how to find the derivative of e^x?

OpenStudy (anonymous):

its e^x

OpenStudy (anonymous):

f(x)=e^x f'(x)=e^x * x'

OpenStudy (anonymous):

replace the x with all that messy stuff and find derivative

OpenStudy (anonymous):

why isnt the f'(x) = to -4e^-4x^2

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

f(x)=(e)^(-4x^(2)) f'(x)=(e)^(-4x^(2)) * (-4x^(2))' f'(x)=(e)^(-4x^(2)) * (2*-4x^(2-1)) f'(x)=(e)^(-4x^(2)) * (-8x^(1)) f'(x)=(e)^(-4x^(2)) * (-8x)

OpenStudy (anonymous):

for f''(x) you use the product rule

OpenStudy (anonymous):

can u please show what the answer is? im having trouble with these

OpenStudy (anonymous):

the e is what is messing me up mainly... also all the powers

OpenStudy (ash2326):

We have \[f(x)=e^{-4x^2}\] Let's find f'(x) \[f'(x)= \frac{d}{dx} e^{-4x^2}\] we get \[f'(x)= e^{-4x^2}\frac{d}{dx} (-4x^2)\] so we get \[f'(x)= e^{-4x^2}(-8x)\]

OpenStudy (anonymous):

understood

OpenStudy (ash2326):

Now let's find f''(x) or second derivative of x \[f''(x)=\frac{d}{dx}( e^{-4x^2}(-8x))\] We'll use product rule here \[(fg)'=f'g+g'f\] so we get here \[f''(x)=(-8x)\frac{d}{dx} (e^{-4x^2}) +e^{-4x^2}\frac{d}{dx} (-8x)\] we get now \[f''(x)=(-8x)(e^{-4x^2}\frac{d}{dx}(-4x^2) ) +e^{-4x^2}(-8)\] so we get \[f''(x)=(-8x)(e^{-4x^2}(-8x) ) +e^{-4x^2}(-8)\] let's simplify now \[f''(x)=(-8x)\times (-8x)e^{-4x^2} -8e^{-4x^2}\] we get finally \[f''(x)=64x^2e^{-4x^2} -8e^{-4x^2}\]

OpenStudy (ash2326):

@mariomintchev Did you understand?

OpenStudy (anonymous):

im soaking it all in right now. im gonna write it all out on paper.

OpenStudy (ash2326):

Yeah , call me if you need help:D

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

so the derivative of e raised to some power is that and the derivative of the power?

OpenStudy (ash2326):

yeah \[\frac{d}{dx} e^x= e^x \time \frac{d}{dx} x\]

OpenStudy (ash2326):

sorry \[\frac{d}{dx} e^x=e^x \times \frac{d}{dx} (x)\]

OpenStudy (anonymous):

you're fine. i understood.

OpenStudy (ash2326):

Great :D

OpenStudy (anonymous):

how would i enter this into a hw site... it keeps telling me that im making a syntax error

OpenStudy (anonymous):

the answer to f"(x)

OpenStudy (ash2326):

Maybe like f''(x)=64x^(2)e^(-4x^(2))-8e^(-4x^2) or e^(-4x^(2))(64x^(2)-8)

OpenStudy (anonymous):

ok thanks! :D